A ball is kicked such that it leaves the ground at an upwards angle of 35° above the horizontal with an initial velocity of 17 m s−1.

How far away does the ball hit the ground? (in m to 2.s.f)

(note: ignore the effects of air resistance when solving this problem)
1
Expert's answer
2019-03-28T05:04:06-0400

Assume accelerations is "g=9.8\,m\,s^{-2}" , angle between initial velocity and horizontal line "\theta=35^ \circ"

The distance travelled:


"d=\frac{v^2 \sin{2\theta}}{g}=\frac{17^2 \sin{70^\circ}}{9.8} m = 27.7 m"